OpenStudy (anonymous):

A motorboat takes 5 hours to travel 200 km going upstream. The return trip takes 4 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (perl):

we can use the equation: distance = rate * time

OpenStudy (anonymous):

Let me try. 200 = r x 5 5/200 = 40 r=40 (Slower) then 40-1 = 39 (Faster )

OpenStudy (anonymous):

40+1 = 41

OpenStudy (anonymous):

Still water rate?

OpenStudy (perl):

Let r = rate of boat in still water c = rate of river current For upstream we have : 200 = ( r - c) * 5 Downstream we have : 200 = ( r + c) * 4

OpenStudy (anonymous):

200 = (40-c)*5 (Up stream) 200= (40 + c)*5 (Downstream)

OpenStudy (perl):

r is unknown

OpenStudy (anonymous):

I thought R -= 40 from my 1st reply

OpenStudy (anonymous):

R= 40 Let me try. 200 = r x 5 5/200 = 40 r=40 (Slower)

OpenStudy (perl):

see there are actually two unknowns here the rate of the boat in still water, and the rate of the river stream. your equation does not reflect this .

OpenStudy (anonymous):

I am so lost..

OpenStudy (perl):

go over the equation in my last post above.

OpenStudy (anonymous):

ok .. Let r = rate of boat in still water c = rate of river current For upstream we have : 200 = ( r - c) * 5 Downstream we have : 200 = ( r + c) * 4 How can I do any more if both r and c are unknown

OpenStudy (perl):

we have two equations upstream 200 = ( r - c) * 5 downstream : 200 = ( r + c) * 4

OpenStudy (perl):

we can solve a system of two equations in two unknowns

OpenStudy (anonymous):

Up = 40 Down = 50

OpenStudy (anonymous):

Not sure I see how that answers rates?

OpenStudy (perl):

They have a step by step solution there. :)

OpenStudy (anonymous):

Let r = rate of boat in still water c = rate of river current For upstream we have : 200 = (45 - 5) * 5 Downstream we have : 200 = ( 45 + 5) * 4

OpenStudy (anonymous):

Thank you!!!